A rock is thrown up from the ground at a velocity of 84 feet per second. The formula h = –16t^2 + 84t gives the rock’s height in feet after t seconds. What is the maximum height of the rock?
Answer by stanbon(75887)

A rock is thrown up from the ground at a velocity of 84 feet per second. The formula h = –16t^2 + 84t gives the rock’s height in feet after t seconds. What is the maximum height of the rock?
-----
max occurs when x = -b/(2a) = -84/(2*-16) = 2.6250 seconds
-----------------------
f(2.6250) = -16(2.6250^2) + 84(2.6250) = 110.25 ft
------
Cheers,
Stan H.